# ImagenHub: Standardizing the evaluation of conditional image generation models

### v1

@@ 1 Introduction
With the recent development of diffusion models, image generation has quickly become one of the most popular research areas in AI. To enable controllability in the image generation process, a myriad of conditional image generation models have been proposed in the past year. Diverse set of conditions have been attempted to steer the diffusion process. One of the most popular task is text-guided image generation (Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022), which aims to ground on a text prompt to generate the corresponding image. Besides that, there are also subject-conditioned image generation (Gal et al., 2022; Ruiz et al., 2023), text-guided image editing (Brooks et al., 2023), multi-subject-conditioned image generation (Kumari et al., 2023), style-guided image generation (Sohn et al., 2023), etc. These different tasks aim to serve different types of downstream applications by enabling subject-level, background-level, style-level controls. The field is evolving at an unprecedented pace and lots of improvements have been reported in the published papers. However, one glaring issue we observed is the published work are highly inconsistent in their experimental setups. To summarize, the inconsistencies mainly come from three aspects, namely dataset, inference and evaluation:
@@
  • We found that evaluation results from the published papers from are generally consistent with our evaluation. 83% of the published result ranking is consistent with our ranking.
@@ 3 Related Work
AI-generated Image Assessment. Evaluating AI-generated images holistically is a complex and open problem  (Salimans et al., 2016). Researchers have proposed various automatic metrics. In the image quality aspect, Inception score  (Salimans et al., 2016), FID (Heusel et al., 2017) are often used. These methods rely on the statistics on an InceptionNet pre-trained on the ImageNet dataset. Despite the fact that they are widely adapted due to the sensitivity to small changes in the image, those metrics are not an ideal solution as they are biased towards the ImageNet dataset and result in inadequate evaluation  (Borji, 2021). Later works like LPIPS  (Zhang et al., 2018) and DreamSim  (Fu et al., 2023) proposed better ways to measure the perceptual similarity. In the semantic consistency aspect, the CLIP score  (Hessel et al., 2021) is often used to measure the vision-language alignment between the generated image and the prompt. Researchers also worked on alternative methods such as BLIP score  (Li et al., 2022) and ImageReward  (Xu et al., 2023). However, in some tasks like subject-driven image generation and editing, the automatic measurement of semantic consistency is still an open problem. One long-established yet effective approach to assessing AI-generated image performance is to rely on human annotators to assess the visual quality  (Denton et al., 2015; Isola et al., 2017; Meng et al., 2021; Chen et al., 2023). The downside is that it entails a reliance on human judgment, which can introduce subjectivity and potentially limit scalability. To mitigate the downsides, the human evaluation design has to be unambiguous and easy to follow.
@@ 4.1 Human Evaluation Metrics
where $g$ is a modularized function to compute the consistency between $y$ and a single condition $c$. We set $g(y,c_{1})\in[0,0.5,1]$, where 0 means inconsistent, 0.5 means partially consistent and 1 means totally consistent. With this formulation, as long as the output is inconsistent with any of the conditions, the evaluation score should become zero. Otherwise, the aggregation function will pick the lowest consistency score from all the conditions. On the other hand, perceptive quality measures the image quality, i.e. whether the image contains artifacts, is blurry, or has an unnatural sense. we set perceptive quality $PQ\in[0,0.5,1]$, where 0 means extremely poor quality, 0.5 means the image has an acceptable quality and 1 means high quality. In these experiments, each model is rated individually. We train human raters to estimate the $g$ function and $PQ$ function with comprehensive guidelines in  Table 1. We derive $O=\sqrt{SC\times PQ}$ as the overall rating of a model. One benefit of using geometric mean as the design choice is that the rating is penalized when one of the aspect scores is too low. We further studied the configuration in section 5.
@@
Recognizable Litte/None Little/None 1.0
@@
Figure 4: Model performance and variance in each subtask.
@@ 4.2 Dataset and available models
We present a standardized dataset for each type of task. The information of the datasets is shown in  Table 10. Some models have different standards of inputs in one task. For example, in the text-guided image editing task, DiffEdit is global description-guided while InstructPix2Pix is instruction-guided. We manually created the equivalent meaning prompts for both methods so they can be aligned. All datasets contain a huge variety of test cases to mimic the diversity in real-life situations. We hosted all of our datasets on HuggingFace dataset for easy access and maintenance. Here we demonstrate all the evaluated models in Table 3.
@@ 5 Experimental Results
Experiment Setup. All the models either used the default setting from the official implementation or the setting suggested in HuggingFace documentation  (von Platen et al., 2022). We disabled negative prompts and any prompt engineering tricks to ensure a fair comparison. We conducted human evaluation by recruiting participants from Prolific and Amazon Mechanical Turk to rate the images, and our own researchers also took part in the image rating process. We assigned 3 raters for each model and computed the SC score, PQ score, and Overall human score. Then we computed the Fleiss kappa  (Fleiss & Cohen, 1973) for each mentioned score. We also computed Krippendorff’s Alpha  (Krippendorff, 2011), which is expected to yield a higher value than Fleiss kappa. This distinction arises from the nature of the rating categories, with Fleiss’ Kappa assuming nominal categories and Krippendorff’s Alpha accommodating ordinal inputs. Both Fleiss’ Kappa and Krippendorff’s Alpha are bounded within the range of [-1, 1], where the value $>0$ indicates an agreement and closer proximity to 1 indicates a higher degree of agreement.
Results. In Figure 4, we present an overview of the model performance across various tasks. Our findings indicate that the performance of the current models is generally underwhelming, with the exception being Text-guided Image Generation and Subject-driven Image Generation, which has models reaching higher than 0.6 on both SC and PQ average. The detailed report on each model’s performance is shown on  Table 4. We noticed that the overall automated metrics’ correlation with SC score and PQ score in each task is below 0.2 except subject-driven image editing task. Metric values are at  Table 7 and  Table 8.
@@
SDXL-Inpainting 0.14 0.2729 0.49$\pm$0.05 0.51$\pm$0.02 0.37$\pm$0.05 0.50 0.72
SD-Inpainting 0.20 0.2676 0.28$\pm$0.04 0.27$\pm$0.10 0.17$\pm$0.07 0.31 0.49
GLIDE 0.21 0.2578 0.20$\pm$0.05 0.48$\pm$0.06 0.16$\pm$0.05 0.33 0.56
BlendedDiffusion 0.34 0.2594 0.12$\pm$0.03 0.11$\pm$0.03 0.05$\pm$0.02 0.36 0.44
@@
MagicBrush 0.19 0.2675 0.51$\pm$0.01 0.65$\pm$0.06 0.47$\pm$0.02 0.44 0.67
InstructPix2Pix 0.27 0.2616 0.29$\pm$0.01 0.70$\pm$0.06 0.27$\pm$0.02 0.55 0.74
Prompt-to-prompt 0.36 0.2674 0.17$\pm$0.05 0.55$\pm$0.09 0.15$\pm$0.06 0.36 0.53
CycleDiffusion 0.21 0.2692 0.17$\pm$0.03 0.56$\pm$0.11 0.14$\pm$0.04 0.41 0.63
SDEdit 0.59 0.2872 0.04$\pm$0.03 0.56$\pm$0.12 0.04$\pm$0.03 0.13 0.13
Text2Live 0.06 0.2628 0.02$\pm$0.01 0.82$\pm$0.04 0.02$\pm$0.02 0.10 0.17
DiffEdit 0.16 0.2425 0.02$\pm$0.01 0.23$\pm$0.04 0.01$\pm$0.01 0.24 0.24
Pix2PixZero 0.57 0.2510 0.01$\pm$0.00 0.48$\pm$0.09 0.01$\pm$0.01 0.37 0.37
@@
ControlNet 0.86 0.2555 0.42$\pm$0.05 0.19$\pm$0.04 0.23$\pm$0.04 0.37 0.57
UniControl 0.90 0.2604 0.38$\pm$0.07 0.20$\pm$0.06 0.23$\pm$0.07 0.36 0.58
Table 4: All the evaluated models from seven core tasks. $\text{Fleiss}_{\bar{O}}$ and $\text{Kd}_{\bar{O}}$, denoting Fleiss’ Kappa and Krippendorff’s alpha for the overall average, respectively. We have more automated metric results in Appendix Table 7.
@@ 5.1 Discovery and Insights
Text-guided Image Editing. One key requirement is to edit the image precisely and keep the background untouched. This requirement is indeed challenging because the network has to understand of editing region from the semantic inputs. We discovered that Prompt-to-Prompt, Pix2PixZero, and SDEdit, despite generating high-quality images, often result in completely different backgrounds. We also spotted that in many cases Text2Live will simply return the input as output, this phenomenon also occasionally happened in other models. For paper claims, our evaluation ranking aligns with the findings from CycleDiffsuion, InstructPix2Pix, MagicBrush, Prompt-to-Prompt, and DiffEdit. We found our evaluation ranking does not align with Pix2PixZero, since their paper only tested on word swapping examples, which is not able to generalize to more complex edits.
@@
Multi-concept Image Composition. Our evaluations validates that CustomDiffusion is consistently better than the other two models. However, while it learns the given multiple subjects’ features better, it could fail to follow the prompts in many cases, especially on actions and positional words. In contrast, DreamBooth learns the correct subjects in some cases, and TextualInversion almost never learn the correct subjects. Interestingly, in some cases where DreamBooth does not learn the correct subjects, it could still follow the prompts correctly.
Subject-driven Image Editing. It is essential to modify the subject from the source to the target without causing excess changes to the background. Human evaluation is also conducted in DreamEdit for the comparison between Photoswap and DreamEdit, but our rankings differ due to varying evaluation criteria. PhotoSwap can adapt to the target subject from the source naturally in most of the cases, but rarely preserves the background well. DreamEdit maintains the context in most of the cases but sometimes leaves observable distortions at the edge of contextualization. BLIP-Diffusion fails the adaptation most of the time, compromising for a more realistic generation.
@@ 5.2 Ablation Study
Method of overall human score computation. We set the overall score of the model as the geometric mean of SC and PR score (i.e. $O=\sqrt{SC\times PQ}$). But we also explored the weighted sum setting $O=\alpha\times SC+\beta\times PQ$, where both $\alpha$ and $\beta$ are in [0, 1]. We conducted an experiment and found that the weighted sum setting yields a different ranking in the models. Take the text-guided image editing task as an example, Text2Live outperforms CycleDiffusion in the weighted sum setting even though we found that CycleDiffusion performs better in the human examination. We investigated and found that a majority of results in Text2Live simply return the input as output (in that case SC=0 and PQ=1). We tried adjusting the weightings but it still failed to reflect the actual performance of the model. Thus we decided to use the geometric mean setting to penalize the models.
@@
Design choice of human evaluation metric range. When it comes to human evaluation on a massive scale, it’s essential to create a system that’s easy to understand and quick to use. In this investigation, we undertake an exploration into how different settings in this evaluation method can affect the results. Initially, our approach entailed the utilization of a range encompassing [0, 0.5, 1, 2] for both the Semantic Consistency (SC) score and the Perception Quality (PQ) score, where 2 means the image is perfect. However, this configuration yielded suboptimal results in the Fleiss Kappa. Subsequently, an alternative configuration was employed, narrowing the range to [0, 1] for both the SC and PQ scores. This adjustment, while accommodating a binary classification, was observed to yield values that were overly polarized and extreme in nature. To find the right balance between keeping values in a reasonable range and making sure the evaluation method is reliable, our resolution was to define a range of [0, 0.5, 1] while providing explicit and unambiguous guidelines.
@@ Conclusion
In this paper, we propose ImagenHub as a continuous effort to unify all the efforts in the conditional image generation into a library to ease the access of these models. We standardize the dataset and evaluation of these models to build our ImagenHub Leaderboard. We hope this leaderboard can deliver more reproducible and fair environment for researchers to visualize the progress in these field. A limitation in this work is the reliance on human raters, which is not only expensive and time-consuming. In the future, we plan to work on developing more generic automatic evaluation methods to approximate the human ratings to help people develop better models.
@@ A.1 More metrics results
Model LPIPS ↓ DINO ↑ CLIP-I ↑ DreamSim ↓ CLIP ↑
@@
DeepFloydIF N/A N/A N/A N/A 0.2814
Stable Diffusion XL N/A N/A N/A N/A 0.2886
Dalle-2 N/A N/A N/A N/A 0.2712
OpenJourney N/A N/A N/A N/A 0.2814
Stable Diffusion N/A N/A N/A N/A 0.2899
@@
SDXL-Inpainting 0.14 0.88 0.93 0.16 0.2729
SD-Inpainting 0.20 0.84 0.91 0.21 0.2676
GLIDE 0.21 0.90 0.94 0.14 0.2578
BlendedDiffusion 0.34 0.79 0.86 0.27 0.2594
@@
MagicBrush 0.19 0.87 0.91 0.18 0.2675
InstructPix2Pix 0.27 0.79 0.88 0.26 0.2616
Prompt-to-prompt 0.36 0.68 0.80 0.36 0.2674
CycleDiffusion 0.21 0.82 0.89 0.22 0.2692
SDEdit 0.59 0.55 0.74 0.51 0.2872
Text2Live 0.07 0.95 0.96 0.07 0.2628
DiffEdit 0.16 0.82 0.86 0.26 0.2425
Pix2PixZero 0.57 0.53 0.76 0.50 0.2510
@@
SuTI 0.77 0.69 0.81 0.32 0.2895
DreamBooth 0.77 0.61 0.80 0.40 0.2847
BLIP-Diffusion 0.77 0.63 0.82 0.36 0.2729
TextualInversion 0.81 0.37 0.65 0.61 0.2680
DreamBooth-Lora 0.82 0.36 0.70 0.61 0.2988
@@
PhotoSwap 0.34 0.65 0.83 0.37 0.2846
DreamEdit 0.22 0.74 0.87 0.28 0.2855
BLIP-Diffusion 0.25 0.80 0.89 0.23 0.2901
@@
CustomDiffusion 0.79 0.50 0.70 0.55 0.2929
DreamBooth 0.78 0.39 0.64 0.66 0.2993
TextualInversion 0.80 0.44 0.66 0.65 0.2548
@@
ControlNet 0.86 0.29 0.67 0.68 0.2555
UniControl 0.90 0.29 0.70 0.69 0.2604
Table 7: All automated metrics for the evaluated models.

### v3

@@ 1 Introduction
With the recent development of diffusion models, image generation has quickly become one of the most popular research areas in AI. To enable controllability in the image generation process, a myriad of conditional image generation models have been proposed in the past year. Diverse set of conditions have been attempted to steer the diffusion process. One of the most popular tasks is text-guided image generation (Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022), which aims to ground on a text prompt to generate the corresponding image. Besides that, there are also subject-conditioned image generation (Gal et al., 2022; Ruiz et al., 2023), text-guided image editing (Brooks et al., 2023), multi-subject-conditioned image generation (Kumari et al., 2023), style-guided image generation (Sohn et al., 2023), etc. These different tasks aim to serve different types of downstream applications by enabling subject-level, background-level, style-level controls. The field is evolving at an unprecedented pace and lots of improvements have been reported in the published papers. However, one glaring issue we observed is the published work are highly inconsistent in their experimental setups. To summarize, the inconsistencies mainly come from three aspects, namely dataset, inference and evaluation:
@@
  • We found that evaluation results from the published papers are generally consistent with our evaluation. 83% of the published result ranking is consistent with our ranking.
@@ 3 Related Work
AI-generated Image Assessment. Evaluating AI-generated images holistically is a complex and open problem  (Salimans et al., 2016). Researchers have proposed various automatic metrics. In the image quality aspect, Inception score  (Salimans et al., 2016), FID (Heusel et al., 2017) are often used. These methods rely on statistics from an InceptionNet pre-trained on the ImageNet dataset. Despite being widely adopted due to their sensitivity to small changes in images, these metrics are not ideal. They are biased towards the ImageNet dataset, resulting in inadequate evaluations  (Borji, 2021). Later works like LPIPS  (Zhang et al., 2018) and DreamSim  (Fu et al., 2023) proposed better ways to measure the perceptual similarity. In the semantic consistency aspect, the CLIP score  (Hessel et al., 2021) is often used to measure the vision-language alignment between the generated image and the prompt. Researchers also worked on alternative methods such as BLIP score  (Li et al., 2022) and ImageReward  (Xu et al., 2023). However, in some tasks like subject-driven image generation and editing, the automatic measurement of semantic consistency is still an open problem. One long-established yet effective approach to assessing AI-generated image performance is to rely on human annotators to assess the visual quality  (Denton et al., 2015; Isola et al., 2017; Meng et al., 2021; Chen et al., 2023). The downside is that it entails a reliance on human judgment, which can introduce subjectivity and potentially limit scalability. To mitigate the downsides, the human evaluation design has to be unambiguous and easy to follow.
@@ 4.1 Human Evaluation Metrics
where $g$ is a modularized function to compute the consistency between $y$ and a single condition $c$. We set $g(y,c_{1})\in[0,0.5,1]$, where 0 means inconsistent, 0.5 means partially consistent and 1 means fully consistent. With this formulation, as long as the output is inconsistent with any of the conditions, the evaluation score should become zero. Otherwise, the aggregation function will pick the lowest consistency score from all the conditions. On the other hand, perceptive quality measures the image quality, i.e. whether the image contains artifacts, is blurry, or has an unnatural sense. we set perceptive quality $PQ\in[0,0.5,1]$, where 0 means extremely poor quality, 0.5 means the image has an acceptable quality and 1 means high quality. In these experiments, each model is rated individually. We train human raters to estimate the $g$ function and $PQ$ function with comprehensive guidelines in  Table 1. We derive $O=\sqrt{SC\times PQ}$ as the overall rating of a model. One benefit of using geometric mean as the design choice is that the rating is penalized when one of the aspect scores is too low. We further studied the configuration in section 5.
@@
Recognizable Little/None Little/None 1.0
@@
Figure 4: Model performance and standard deviation in each task.
@@ 4.2 Dataset and available models
We present a standardized dataset for each type of task. The information of the datasets is shown in  LABEL:tab:dataset. Some models have different standards of inputs in one task. For example, in the text-guided image editing task, DiffEdit is global description-guided while InstructPix2Pix is instruction-guided. We manually created the equivalent meaning prompts for both methods so they can be aligned. All datasets contain a huge variety of test cases to mimic the diversity in real-life situations. We hosted all of our datasets on the HuggingFace dataset for easy access and maintenance. Here we demonstrate all the evaluated models in Table 3.
@@ 5 Experimental Results
Experiment Setup. All the models either used the default setting from the official implementation or the setting suggested in HuggingFace documentation  (von Platen et al., 2022). We disabled negative prompts and any prompt engineering tricks to ensure a fair comparison. We conducted human evaluation by recruiting participants from Prolific to rate the images, and our own researchers also took part in the image rating process. We assigned 3 raters for each model and computed the SC score, PQ score, and Overall human score. Then we computed the Fleiss kappa  (Fleiss & Cohen, 1973) for each mentioned score. We also computed Krippendorff’s Alpha  (Krippendorff, 2011), which is expected to yield a higher value than Fleiss kappa. This distinction arises from the nature of the rating categories, with Fleiss’ Kappa assuming nominal categories and Krippendorff’s Alpha accommodating ordinal inputs. Both Fleiss’ Kappa and Krippendorff’s Alpha are bounded within the range of [-1, 1], where the value $>0$ indicates an agreement and closer proximity to 1 indicates a higher degree of agreement.
Results. In Figure 4, we present an overview of the model performance across various tasks. Our findings indicate that the performance of the current models is generally underwhelming, with the exception being Text-guided Image Generation and Subject-driven Image Generation, which have models reaching higher than 0.6 on both SC and PQ averages. The detailed report on each model’s performance is shown on  Table 4. We noticed that the overall automated metrics’ correlation with the SC score and PQ score in each task is below 0.2 except subject-driven image editing task. Metric values are at  Table 7 and  Table 8.
@@
SDXL-Inpainting 0.15 0.2729 0.49$\pm$0.05 0.51$\pm$0.02 0.37$\pm$0.05 0.50 0.72
SD-Inpainting 0.21 0.2676 0.28$\pm$0.04 0.27$\pm$0.10 0.17$\pm$0.07 0.31 0.49
GLIDE 0.18 0.2578 0.20$\pm$0.05 0.48$\pm$0.06 0.16$\pm$0.05 0.33 0.56
BlendedDiffusion 0.33 0.2594 0.12$\pm$0.03 0.11$\pm$0.03 0.05$\pm$0.02 0.36 0.44
@@
MagicBrush 0.22 0.2675 0.51$\pm$0.01 0.65$\pm$0.06 0.47$\pm$0.02 0.44 0.67
InstructPix2Pix 0.32 0.2616 0.29$\pm$0.01 0.70$\pm$0.06 0.27$\pm$0.02 0.55 0.74
Prompt-to-prompt 0.40 0.2674 0.17$\pm$0.05 0.55$\pm$0.09 0.15$\pm$0.06 0.36 0.53
CycleDiffusion 0.28 0.2692 0.17$\pm$0.03 0.56$\pm$0.11 0.14$\pm$0.04 0.41 0.63
SDEdit 0.61 0.2872 0.04$\pm$0.03 0.56$\pm$0.12 0.04$\pm$0.03 0.13 0.13
Text2Live 0.17 0.2628 0.02$\pm$0.01 0.82$\pm$0.04 0.02$\pm$0.02 0.10 0.17
DiffEdit 0.22 0.2425 0.02$\pm$0.01 0.23$\pm$0.04 0.01$\pm$0.01 0.24 0.24
Pix2PixZero 0.60 0.2510 0.01$\pm$0.00 0.48$\pm$0.09 0.01$\pm$0.01 0.37 0.37
@@
ControlNet 0.80 0.2555 0.42$\pm$0.05 0.19$\pm$0.04 0.23$\pm$0.04 0.37 0.57
UniControl 0.82 0.2604 0.38$\pm$0.07 0.20$\pm$0.06 0.23$\pm$0.07 0.36 0.58
Table 4: All the evaluated models from seven core tasks. Overall is the average of all $\sqrt{SC\times PQ}$. $\text{Fleiss}_{\bar{O}}$ and $\text{Kd}_{\bar{O}}$, denoting Fleiss’ Kappa and Krippendorff’s alpha for the overall average, respectively. We have more automated metric results in Appendix Table 7 and correlations in Table 8.
@@ 5.1 Discovery and Insights
Text-guided Image Editing. One key requirement is to edit the image precisely and keep the background untouched. This requirement is indeed challenging because the network has to understand of editing region from the semantic inputs. We discovered that Prompt-to-Prompt, Pix2PixZero, and SDEdit, despite generating high-quality images, often result in completely different backgrounds. We also spotted that in many cases Text2Live will simply return the input as output, this phenomenon also occasionally happened in other models. For paper claims, our evaluation ranking aligns with the findings from CycleDiffsuion, InstructPix2Pix, MagicBrush, Prompt-to-Prompt, and DiffEdit. We found our evaluation ranking does not align with Pix2PixZero, since their paper only tested on word-swapping examples, which is not able to generalize to more complex edits.
@@
Multi-concept Image Composition. Our evaluations validate that CustomDiffusion is consistently better than the other two models. However, while it learns the given multiple subjects’ features better, it could fail to follow the prompts in many cases, especially on actions and positional words. In contrast, DreamBooth learns the correct subjects in some cases, and TextualInversion rarely learns the correct subjects. Interestingly, in some cases where DreamBooth does not learn the correct subjects, it could still follow the prompts correctly.
Subject-driven Image Editing. It is essential to modify the subject from the source to the target without causing excess changes to the background. Human evaluation is also conducted in DreamEdit for the comparison between Photoswap and DreamEdit, but our rankings differ due to varying evaluation criteria. PhotoSwap can adapt to the target subject from the source naturally in most cases but rarely preserves the background well. DreamEdit maintains the context in most cases but sometimes leaves observable distortions at the edge of contextualization. BLIP-Diffusion fails the adaptation most of the time, compromising for a more realistic generation.
@@ 5.2 Ablation Study
Method of overall human score computation. We set the overall score of the model as the geometric mean of SC and PR score (i.e. $O=\sqrt{SC\times PQ}$). But we also explored the weighted sum setting $O=\alpha\times SC+\beta\times PQ$, where both $\alpha$ and $\beta$ are in [0, 1]. We experimented and found that the weighted sum setting yields a different ranking in the models. Take the text-guided image editing task as an example, Text2Live outperforms CycleDiffusion in the weighted sum setting even though we found that CycleDiffusion performs better in the human examination. We investigated and found that a majority of results in Text2Live simply return the input as output (in that case SC=0 and PQ=1). We tried adjusting the weightings but it still failed to reflect the actual performance of the model. Thus we decided to use the geometric mean setting to penalize the models.
@@
Design choice of human evaluation metric range. When it comes to human evaluation on a massive scale, it’s essential to create a system that’s easy to understand and quick to use. In this investigation, we undertake an exploration into how different settings in this evaluation method can affect the results. Initially, our approach entailed the utilization of a range encompassing [0, 0.5, 1, 2] for both the Semantic Consistency (SC) score and the Perception Quality (PQ) score, where 2 means the image is perfect. However, this configuration yielded suboptimal results in the Fleiss Kappa. Subsequently, an alternative configuration was employed, narrowing the range to [0, 1] for both the SC and PQ scores. This adjustment, while accommodating a binary classification, was observed to yield values that were overly polarized and extreme. To find the right balance between keeping values in a reasonable range and making sure the evaluation method is reliable, we resolved to define a range of [0, 0.5, 1] while providing explicit and unambiguous guidelines.
@@ Conclusion
In this paper, we propose ImagenHub as a continuous effort to unify all efforts in conditional image generation into a library, easing access to these models. We standardize the dataset and evaluation of these models to build our ImagenHub Leaderboard. We hope this leaderboard can provide a more reproducible and fair environment for researchers to visualize progress in this field. A limitation of this work is the reliance on human raters, which is expensive and time-consuming. In the future, we plan to develop more generic automatic evaluation methods that approximate human ratings, helping people develop better models.
@@ A.1 More metrics results
Model LPIPS ↓ DINO ↑ CLIP-I ↑ DreamSim ↓ FID ↓ KID ↓ CLIP ↑
@@
DeepFloydIF N/A N/A N/A N/A N/A N/A 0.2814
Stable Diffusion XL N/A N/A N/A N/A N/A N/A 0.2886
Dalle-2 N/A N/A N/A N/A N/A N/A 0.2712
OpenJourney N/A N/A N/A N/A N/A N/A 0.2814
Stable Diffusion N/A N/A N/A N/A N/A N/A 0.2899
@@
SDXL-Inpainting 0.15 0.90 0.94 0.15 59.04 -0.0033 0.2729
SD-Inpainting 0.21 0.85 0.91 0.20 86.23 -0.0008 0.2676
GLIDE 0.18 0.91 0.94 0.14 61.71 -0.0028 0.2578
BlendedDiffusion 0.33 0.80 0.87 0.26 102.73 -0.0012 0.2594
@@
MagicBrush 0.22 0.88 0.92 0.19 77.81 -0.0027 0.2675
InstructPix2Pix 0.32 0.77 0.86 0.30 100.30 -0.0014 0.2616
Prompt-to-prompt 0.40 0.68 0.81 0.36 137.68 0.0003 0.2674
CycleDiffusion 0.28 0.79 0.88 0.27 108.29 -0.0008 0.2692
SDEdit 0.61 0.57 0.76 0.50 156.12 0.0025 0.2872
Text2Live 0.17 0.87 0.92 0.18 68.33 -0.0028 0.2628
DiffEdit 0.22 0.78 0.85 0.31 121.79 0.0069 0.2425
Pix2PixZero 0.60 0.53 0.76 0.52 161.64 0.0072 0.2510
@@
SuTI 0.77 0.69 0.81 0.32 164.97 0.0111 0.2895
DreamBooth 0.77 0.61 0.80 0.40 173.71 0.0057 0.2847
BLIP-Diffusion 0.77 0.63 0.82 0.36 166.53 0.0102 0.2729
TextualInversion 0.81 0.37 0.65 0.61 229.18 0.0122 0.2680
DreamBooth-Lora 0.82 0.36 0.70 0.61 234.95 0.0125 0.2988
@@
PhotoSwap 0.34 0.65 0.83 0.37 130.81 0.0009 0.2846
DreamEdit 0.22 0.74 0.87 0.28 118.95 0.0011 0.2855
BLIP-Diffusion 0.25 0.80 0.89 0.23 102.17 0.0007 0.2901
@@
CustomDiffusion 0.79 0.50 0.70 0.55 207.43 0.0341 0.2929
DreamBooth 0.78 0.39 0.64 0.66 257.60 0.0533 0.2993
TextualInversion 0.80 0.44 0.66 0.65 226.25 0.0491 0.2548
@@
ControlNet 0.80 0.43 0.70 0.56 219.29 0.0330 0.2555
UniControl 0.82 0.42 0.68 0.58 222.57 0.0192 0.2604
Table 7: All automated metrics for the evaluated models based on our benchmark dataset.